OpenStudy (anonymous):

Factor completely: 2x^2 + 9x + 4

jimthompson5910 (jim_thompson5910):

hint: find two numbers that multiply to 8 and add to 9

OpenStudy (anonymous):

2 and 4 + 9 = 17

jimthompson5910 (jim_thompson5910):

not sure what you're saying exactly

jimthompson5910 (jim_thompson5910):

sorry

OpenStudy (anonymous):

2*4=8 + 9=17?

jimthompson5910 (jim_thompson5910):

how about 1 and 8 1*8 = 8 1+8 = 9

jimthompson5910 (jim_thompson5910):

see how I got that?

OpenStudy (anonymous):

ah yeah ok.

jimthompson5910 (jim_thompson5910):

to be honest I wasn't sure what you were doing with adding on 9 to 8 to get 17

OpenStudy (anonymous):

here are my choices: (2x + 2)(x + 2) (2x + 1)(x + 4) (2x + 4)(x + 1) (2x + 2)(x + 4)

jimthompson5910 (jim_thompson5910):

anyways, you break up 9x into 1x + 8x to get 2x^2 + 9x + 4 2x^2 + 1x + 8x + 4 then you factor by grouping

OpenStudy (anonymous):

ok and = 2x^2 + 9x +4

OpenStudy (anonymous):

then it would be 2x+1)(x+4)?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

ok thanks!

jimthompson5910 (jim_thompson5910):

you can check the answer by distributing it back out (or FOILing it back out)